eviant:

Rewrite the equation of the circle (x-2)^2+y^2=3 in general form

eviant:

@dude @Vocaloid @Narad

lowkey:

Do you still need help?

Vocaloid:

 expand (x-2)^2 w/ foil, arrange all the terms in the equation in the order described above

eviant:

@dude x^2-4x+y^2-5=0??

eviant:

@Narad

Narad:

It's the reverse of the 1st question, Complete the squares \[x ^{2}+ax+y ^{2}+by+c=0\] \[x^{2}+ax+a^2/4 +y ^{2}+by+b^2/4=-c+a^2/4+b^2/4\] \[(x+a/2)^2+(y+b/2)^2=-c+a^2/4+b^2/4\] If b doesn't exists, then b=0

Narad:

The -5 is not wright

eviant:

@Vocaloid I got x^2-4x+y^2+1=0?

Narad:

Yes